OpenStudy (anonymous):

Find the sum of the series. The sum from n=0 to infinity of 2^(n+2)/3^n

OpenStudy (mathmath333):

\(\large \color{black}{ \sum_{i=0}^{\infty}\frac{2^{n+2}}{3^n}=\sum_{i=0}^{\infty}\frac{2^n\cdot 2^2}{3^n}\hspace{.33em}\\~\\ =4\sum_{i=0}^{\infty}\frac{2^n}{3^n}\hspace{.33em}\\~\\ \implies \sum_{i=0}^{\infty}\left(\frac{2}{3}\right)^n=\left(\frac{2}{3}\right)^0+\left(\frac{2}{3}\right)^1+\left(\frac{2}{3}\right)^2+\left(\frac{2}{3}\right)^3+\cdots \infty \hspace{.33em}\\~\\ \normalsize \text{let }\sum_{i=0}^{\infty}\frac{2^{n}}{3^n}=x\hspace{.33em}\\~\\ x=\left(\frac{2}{3}\right)^0+\left(\frac{2}{3}\right)^1+\left(\frac{2}{3}\right)^2+\left(\frac{2}{3}\right)^3+\ldots +\infty \hspace{.33em}\\~\\ x=1+\left(\frac{2}{3}\right)\left(1+\left(\frac{2}{3}\right)^1+\left(\frac{2}{3}\right)^2+\ldots \infty\right) \hspace{.33em}\\~\\ x=1+\left(\frac{2}{3}\right)\left(x\right) \hspace{.33em}\\~\\ 3x=\left(3+2x\right) \hspace{.33em}\\~\\ x=3\hspace{.33em}\\~\\ 4\sum_{i=0}^{\infty}\frac{2^{n}}{3^n}=4x=4\times 3=12\hspace{.33em}\\~\\ }\)